Calculate the after-tax cost of the debt under each of the following conditions:

a. rd of 5%, tax rate of 12%

b. rd of 5%, tax rate of 32%

c. rd of 5%, tax rate of 42%
Answer #1

after-tax cost of the debt = pre-tax cost of debt * (1 - tax rate)

a]  

after-tax cost of the debt = 5% * (1 - 12%) = 4.4%

b]

after-tax cost of the debt = 5% * (1 - 32%) = 3.4%

c]

after-tax cost of the debt = 5% * (1 - 42%) = 2.9%